OpenStudy (jasbe):

solve x + y + z = -6 x - y + z = -2 2x + y - 3z = 5

OpenStudy (mathstudent55):

Add the first and second equations, and y will drop out. Then add the second and third equation, and y will drop out again. Use the two equations in two unknowns above to solve for x and z. Then substitute to solve for y.

OpenStudy (jasbe):

@mathstudent55 can you walk me through it please ?

OpenStudy (mathstudent55):

Can you add the first two equations? x + y + z = -6 + x - y + z = -2 -----------------